OpenStudy (anonymous):

Suppose that the amount of algae in a pond doubles every 4 hours. If the pond initially contains 90 pounds of algae, how much algae will be in the pond after 12 hours? 720 pounds 360 pounds 1,440 pounds 114 pounds

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@ACTIONSPORTS90210

OpenStudy (anonymous):

it was A

OpenStudy (anonymous):

it will double 3 times find \(90\times 2^3=90\times 8\)

OpenStudy (anonymous):

sh*t i was to late:(sorry